# Dark matter admixed neutron stars with a realistic nuclear equation of state from chiral nuclear interactions

### v1

@@ IV.2 Proton fraction and direct URCA processes in DANS
$f_{\chi}\ [\%]$ $n(r=0)\ [\mathrm{fm^{-3}}]$ $x_{p}(r=0)$ $r_{OM}^{durca}\ [\mathrm{km}]$
0 0.497 0.182 6.84
3.10 0.539 0.196 6.94
5.41 0.578 0.208 7.08
Table 2: Central baryon number density (second column), central proton fraction (third column) and the radial coordinate $r_{OM}^{durca}$ below which direct URCA processes are possible in a $1.4\ \mathrm{M_{\odot}}$ DANS for different value of the dark matter fraction $f_{\chi}$ (first column). All the results are relative to the case $m_{\chi}=1\ \mathrm{GeV}$.
$f_{\chi}\ [\%]$ $n(r=0)\ [\mathrm{fm^{-3}}]$ $x_{p}(r=0)$ $r_{OM}^{durca}\ [\mathrm{km}]$
0 0.345 0.129 /
3.73 0.377 0.141 2.40
6.06 0.393 0.147 3.13
@@
Table 4: Values for $M^{durca}_{tot}=M_{tot}(n^{durca})$ (second column) and for the maximum gravitational mass $M_{max}$ for DANS sequences for different values of dark matter fraction (first column). The third column reports values of obtained for the corresponding $f_{\chi}$ value. All the results are relative to the case $m_{\chi}=1\ \mathrm{GeV}$.
@@
The calculated values for the threshold baryon number density $n^{durca}$ for having direct URCA processes and the corresponding threshold proton fraction $x_{p}(n^{durca})$ do not depend on the DM fraction, since these two quantities are exclusively determined by the nuclear interactions and particularly are strongly affected by the density dependence of the nuclear symmetry energy Bombaci (2022). In the case of the BL EOS one has $n^{durca}=0.361\ \mathrm{fm^{-3}}$ and $x_{p}(n^{durca})=0.1347$ Bombaci and Logoteta (2018). The stellar total gravitational mass $M_{tot}^{durca}\equiv M_{tot}(n_{c}=n^{durca})$ beyond which direct URCA processes are possible depends instead on the DM fraction. Its value together with the value of the gravitational maximum mass $M_{tot}^{max}$ is reported in Tab. 4 for some values of the DM fraction. As we see the presence of a DM core reduces the value of $M_{tot}^{durca}$ and thus makes possible fast cooling of ”light” neutron stars.
V Constraints from astrophysical observations on DM properties
@@
As discussed in Ivanytskyi et al. (2020), it is possible to give an upper limit to the dark matter particles mass by using the results obtained in figure (7) and comparing them with an estimation of the amount of dark matter that can be accumulated onto a neutron star. Such an estimation requires to evaluate the dark matter accretion rate at every evolutionary stage of the star, from the formation of the progenitor star to the stable neutron star stage.
@@
Dark matter accretion rate onto neutron stars as been discussed in Kouvaris (2008) and the total amount of dark matter accumulated in a given time $t$ can be written as (see Ivanytskyi et al. (2020))
@@
$M_{DM}^{acc}\approx 10^{-14}\Bigl{(}\frac{\rho_{\chi}}{0.3\ \mathrm{GeV/cm^{3}}}\Bigr{)}\Bigl{(}\frac{\sigma_{\chi n}}{10^{-45}\ \mathrm{cm^{2}}}\Bigr{)}\Bigl{(}\frac{t}{1\ \mathrm{Gyr}}\Bigr{)}\ \mathrm{M_{\odot}}$ (9)
@@
Following Ivanytskyi et al. (2020), by using the Einasto profile for $\rho_{\chi}$ Einasto (1965, 1969); Merritt et al. (2006); Baes (2022), assuming $\sigma_{\chi n}\sim 10^{-45}\ cm^{2}$ and $t\sim 10\ \mathrm{Gyr}$, we can estimate $M_{DM}^{acc}\sim(10^{-14}-10^{-13})\ \mathrm{M_{\odot}}$ for a $2\ \mathrm{M_{\odot}}$ mass neutron stard in the disk of our galaxy. This values are to low to put a stringent upper limit on $m_{\chi}$. As discussed in Ivanytskyi et al. (2020), using equation (9), for a $2\ \mathrm{M_{\odot}}$ mass neutron stars in the central region of our galaxy, we find that $f_{\chi}\leq 0.01\%$. Using this estimation, equation (7) give us an upper limit for the dark matter particles:
@@
which is close to the result obtained in Ivanytskyi et al. (2020) ($m_{\chi}\leq 60\ \mathrm{GeV}$) in spite the fact that we are using a different EOS for ordinary matter.
We observe that in the dark matter accretion rate estimation made in Kouvaris (2008), only the dark matter-nucleon interaction cross-section is taken into account. If we assume a non zero self interaction between dark matter particles, the accretion rate and the result in equation (10) may change drastically.
VI Stability of DANS
Figure 10: Critical stability curve for $m_{\chi}=1\ \mathrm{GeV}$. The white dashed line is the $\Delta R=R_{OM}-R_{DM}=0$ contour line.
A solution the two-fluid TOV equations (1) and (2) represents an equilibrium configuration. In this section we present results regarding the stability with respect to small radial perturbation of DANS. We assume that these perturbations don’t modify the $\beta$-equilibrium condition of ordinary matter.
In the one-fluid limit (stars made only of ordinary matter or only of dark matter) it is well known that the critical point is the maximum of the $M(\rho_{c,j})$ curve Shapiro and Teukolsky (1983). Using the EOSs considered in this work we get the following value for the central mass densities of the critical configuration
@@ VI Stability of DANS
In the two-fluids system case we get a critical stability curve in the $(\rho_{cOM},\rho_{cDM})$ plane.
The stability analysis can be done in two different ways. The first method, discussed by Kain in Kain (2020), requires to solve a set of pulsation equations whose solution give the squared radial oscillation frequency $\omega^{2}$. If the squared radial oscillation frequency for the fundamental mode $\omega_{0}^{2}$ is positive, than the corresponding solution is stable while, if $\omega_{0}^{2}<0$, the solution is unstable. This method requires to take the derivative of the EOSs of the fluids which is a problem if the EOSs aren’t smooth enough.
The second method, developed by Henriques, Liddle and Moorhouse in Henriques et al. (1990), allow us to determine the critical stability curve without solving the pulsation equations. Their analysis give the following equation for the critical stability curve
@@
where $M_{tot}$, $N_{i}$ ($i=OM,DM$) are the total gravitational mass and total fluid number of a static solution and $\vec{\sigma}$ is a vector in the parameter space that is simultaneously tangent to the level curves of $M_{tot}$ and $N_{i}$. It can be shown Henriques et al. (1990) that if two of the quantities in (11) are zero, then the third is also zero so, to determine the critical stability curve, we can study the stationary point of $M_{tot}$ along level curve of $N_{DM}$. As observed in Henriques et al. (1990), the stability region in the $(\rho_{cOM},\rho_{cDM})$ plane must contain the stable structure in the one-fluid limit.
@@
Figure (10) shows results obtained for $m_{\chi}=1\ \mathrm{GeV}$. This results are similar to those obtained by Kain in Kain (2021) and show us that we can have stable DANS with both a core and a dark matter halo. We also note that there is a region inside the stability curve where DANS have an ordinary matter central density and/or a dark matter central density that exceed the critical value in the one-fluid limit.
VII Summary

### v2

@@ 4.2 Proton fraction and direct URCA processes in DANS
As is clearly seen in Fig. 8, the presence a DM core produces a significant increase in the proton fraction (particularly in the central region of the star) and small increase in the value of $r_{OM}^{durca}$ Notice that in the case of an ordinary neutron star ($f_{\chi}=0$) with mass $M=1.4\ \mathrm{M_{\odot}}$ described by the BL EOS, direct URCA processes are already possible (Bombaci and Logoteta, 2018). Complementary to Fig. 8, we report in Tab. 2 some of the properties of a DANS with $M_{tot}=1.4\ \mathrm{M_{\odot}}$ in the case $m_{\chi}=1\ \mathrm{GeV}$.
@@
$f_{\chi}$ $\rho_{cOM}$ $n(r=0)$ $x_{p}(r=0)$ $r_{OM}^{durca}$
[…]
0 0.910 0.497 0.182 6.84
3.10 1.000 0.539 0.196 6.94
5.41 1.086 0.578 0.208 7.08
Table 2: Central mass and baryon number density (first and second column), central proton fraction (third column) and the radial coordinate $r_{OM}^{durca}$ below which direct URCA processes are possible in a $1.4\ \mathrm{M_{\odot}}$ DANS for different value of the dark matter fraction $f_{\chi}$ (first column). All the results are relative to the case $m_{\chi}=1\ \mathrm{GeV}$.
$f_{\chi}$ $\rho_{cOM}$ $n(r=0)$ $x_{p}(r=0)$ $r_{OM}^{durca}$
[…]
0 6.076 0.345 0.129 /
3.73 6.688 0.377 0.141 2.40
6.06 6.993 0.393 0.147 3.13
@@
Table 4: Values for $M^{durca}_{tot}=M_{tot}(n^{durca})$ (second column) and for the maximum gravitational mass $M_{max}(f_{\chi})$ (second column) for DANS sequences with different values of dark matter fraction $f_{\chi}$. The calculated values for the threshold baryon number density for having direct URCA processes is $n^{durca}=0.361\ \mathrm{fm^{-3}}$. All the results are relative to the case $m_{\chi}=1\ \mathrm{GeV}$.
@@
The calculated values for the threshold baryon number density $n^{durca}$ for having direct URCA processes and the corresponding threshold proton fraction $x_{p}(n^{durca})$ do not depend on the DM fraction, since these two quantities are exclusively determined by the nuclear interactions and particularly are strongly affected by the density dependence of the nuclear symmetry energy (Bombaci and Lombardo, 1991; Bombaci, 2022). In the case of the BL EOS one has $n^{durca}=0.361\ \mathrm{fm^{-3}}$ and $x_{p}(n^{durca})=0.1347$ (Bombaci and Logoteta, 2018). The stellar total gravitational mass $M_{tot}^{durca}\equiv M_{tot}(n_{c}=n^{durca})$ beyond which direct URCA processes are possible depends instead on the DM fraction. Its value together with the value of the gravitational maximum mass $M_{tot}^{max}$ is reported in Tab. 4 for some values of the DM fraction. As we see the presence of a DM core reduces the value of $M_{tot}^{durca}$ and thus makes possible fast cooling of ”light” neutron stars. Our results are in general agreement with what reported in the recent papers (Ávila et al., 2024; Giangrandi et al., 2024). For example, we find that a DM fraction $f_{\chi}\sim 3\%$ produces a decrease in the critical mass $M_{tot}^{durca}$ for having direct URCA process of about $0.09\,M_{\odot}$, while the authors of reference (Ávila et al., 2024; Giangrandi et al., 2024) find a decrease between 0.10 and 0.13 $M_{\odot}$ for the two cases of ordinary matter EOS they used. However, we find a significant difference when we compare our calculated values for $M_{tot}^{durca}(f_{\chi})$ with those reported in (Ávila et al., 2024; Giangrandi et al., 2024). For example, we predict that a DM fraction $f_{\chi}=3.1\%$ could produce rapid cooling of neutron stars with a quite low mass ($M_{tot}\geq 0.87\,M_{\odot}$) whereas in (Ávila et al., 2024; Giangrandi et al., 2024), it is found $M_{tot}^{durca}=1.80\,M_{\odot}$ for the same DM fraction and in the case of the phenomenological Induced Surface Tension (IST) EOS model for ordinary matter.
We are able to trace the origin of these substantial differences between our results and those of (Ávila et al., 2024; Giangrandi et al., 2024). In fact, large differences in the calculated threshold stellar mass $M_{tot}^{durca}$ for having direct URCA processes are already present when we compare our results (see also Table 4 in (Bombaci and Logoteta, 2018)) with those of Ref. (Ávila et al., 2024; Giangrandi et al., 2024) in the case of ordinary neutron stars ($f_{\chi}=0$). Definitely, these differences are due to the use of distinct EOS models for describing ordinary matter and, as we argue below, they are due to a different density dependence of the nuclear symmetry energy in the considered nuclear matter EOS models.
As already discusses in Section 3.1, the value of the proton fraction $x_{p}$ in $\beta$-stable nuclear matter (and the threshold baryon number density $n^{durca}$ for having direct URCA process) is highly related to the density dependence of the symmetry energy $E_{sym}(n)$. A stiff symmetry energy, as suggested by the PREX-2 measurement (Adhikari et al., 2021) of the neutron skin thickness of the 208Pb nucleus, favors a large proton fraction and a low value ($M_{tot}^{durca}\lesssim 1.45\,M_{\odot}$) for the threshold stellar mass for direct URCA processes (see Fig. 3 in (Reed et al., 2021)), in agreement with our results for ordinary ($f_{\chi}=0$) neutron stars.
In the case of the IST EOS model, used in (Ávila et al., 2024; Giangrandi et al., 2024), the symmetry energy $E_{sym}(n)$ is softer than the one for the BL model and as a consequence the proton fraction $x_{p}(n)$ in $\beta$-stable matter in the case of the IST EOS is much lower than the one of the BL EOS (compare Fig. 4 in (Ávila et al., 2024) with Fig. 3 in (Bombaci and Logoteta, 2018)). Note, also, that for the case of the IST and FSU2R EOS models, used in (Ávila et al., 2024; Giangrandi et al., 2024), the symmetry energy $E_{sym}(n_{0})$ and its slope $L$, at nuclear saturation density, are in disagreement with the values derived by the measured neutron skin thickness of the 208Pb nucleus by the PREX-2 experiment. In addition these two physical quantities violate the bounds imposed by the unitary Fermi gas limit (Tews et al., 2017) in the case of the IST EOS.
Besides, in calculating the proton fraction in $\beta$-stable nuclear matter the authors of references (Ávila et al., 2024; Giangrandi et al., 2024) do not include negative muons ($\mu^{-}$). In the BL EOS model muons appear at a baryon density $n=0.123\ \mathrm{fm^{-3}}$ (i.e. below the nuclear matter saturation density). Any reliable calculation of the proton fraction in $\beta$-stable nuclear matter requires the presence of muons.
Therefore, our results on the effect of dark matter on the proton fraction in DANS and the possibility of having fast cooling already in stars with a mass below one solar mass are very interesting and different from what has been reported in the literature so far.
5 Constraints from astrophysical observations on DM properties
@@
As discussed in (Ivanytskyi et al., 2020), it is possible to give an upper limit to the mass $m_{\chi}$ of the dark matter particles by using the results reported in Fig. (7) and comparing them with the estimate of the amount of dark matter that can be accumulated onto a neutron star. To make such an estimate, it is necessary to evaluate the dark matter accretion rate at each evolutionary stage of the star, from the formation of the progenitor star up to the stable neutron star phase.
The accretion of DM on massive stars which form neutron stars, has been most carefully investigated during their main sequence phase (Scott et al., 2009; Lopes et al., 2019) and during the stable neutron star phase (Kouvaris, 2008; Kouvaris and Tinyakov, 2011).
@@
Dark matter accretion rate onto neutron stars has been discussed in (Kouvaris, 2008) and the total amount of dark matter accumulated onto a $2\,\mathrm{M_{\odot}}$ neutron star in a given time $t$ (i.e. a neutron star with age $t$) can be estimated (Del Popolo et al., 2020; Ivanytskyi et al., 2020) as
@@
$\begin{split}M_{DM}^{acc}\approx 10^{-14}&\Bigl{(}\frac{\rho_{\chi}}{0.3\ \mathrm{GeV/cm^{3}}}\Bigr{)}\times\\ &\times\Bigl{(}\frac{\sigma_{\chi n}}{10^{-45}\ \mathrm{cm^{2}}}\Bigr{)}\Bigl{(}\frac{t}{1\ \mathrm{Gyr}}\Bigr{)}\ \mathrm{M_{\odot}}\end{split}$ (9)
@@
To model the dark matter density profile $\rho_{\chi}$, we can use the Einasto profile (Einasto, 1965, 1969), since it is in good agreement with numerical simulations and data from the rotation curve of the Milky Way (Sofue, 2013; Pato et al., 2015; Einasto, 1965, 1969; Merritt et al., 2006; Baes, 2022). As a function of the distance $r$ from the galactic center we can write
$\frac{\rho_{\chi}(r;\alpha,r_{-2},\rho_{-2})}{\rho_{-2}}=\exp\Biggl{\{}-\frac{2}{\alpha}\Biggl{[}\Bigl{(}\frac{r}{r_{-2}}\Bigr{)}^{\alpha}-1\Biggr{]}\Biggr{\}}$
where $\alpha$ is a parameter that determine the curvature of the profile, $\rho_{-2}$ is a normalization constant and $r_{-2}$ is defined by the following condition:
$\frac{d\ln\rho_{\chi}(r)}{d\ln r}\Biggr{|}_{r=r_{-2}}=-2$
As discussed in (Del Popolo et al., 2020), we can set $r_{-2}=15\ \mathrm{kpc}$ and fix $\rho_{-2}$ by using the fact that in the solar neighborhood ($r=R_{\odot}=8\ \mathrm{kpc}$) we should have $\rho_{\chi}(R_{\odot})=0.420\ \mathrm{GeV/cm^{3}}$ (Pato et al., 2015). Depending on the value of the $\alpha$ parameter, we get different density profiles. The total mass of accreted DM onto the neutron star is not very sensitive to the value of the $\alpha$ parameter in the galaxy disc. For a typical neutron star in the solar neighborhood one can get $M_{DM}^{acc}\approx 10^{-11}\ \mathrm{M_{\odot}}$ (Del Popolo et al., 2020), while the same estimate for PSR J0348+0432 and PSR J0740+6620 leads to $M_{DM}^{acc}\approx 10^{-13}-10^{-14}\ \mathrm{M_{\odot}}$ (Ivanytskyi et al., 2020) depending on the value of the parameters. These values for $M_{DM}^{acc}$ are too low to produce any significant effect on the neutron star structure and to put a stringent upper limit on $m_{\chi}$.
To get an upper limit for the mass of the DM particles, we should consider a very old neutron star ($t\sim 10\ \mathrm{Gyr}$) in a DM rich region like the galactic center. At a distance $r\sim 10^{-5}\ \mathrm{pc}$, we can have up to $M_{DM}^{acc}\approx 10^{-4}-10^{-3}\ \mathrm{M_{\odot}}$ depending on the value of the parameter $\alpha$ in the Einasto profile (Del Popolo et al., 2020). In (Ivanytskyi et al., 2020), using this approach and taking also into account the DM accretion during the main sequence phase of the neutron star progenitor, a value of $f_{\chi}=M_{DM}^{acc}/M_{tot}\approx 0.01\%$ is obtained for a $2\ \mathrm{M_{\odot}}$ neutron star in the galactic center.
The latter estimate for $f_{\chi}$ is represented by the blue dashed line in Fig. (7). Combining this latter estimate with our calculations reported in Fig. (7), our equation (7) gives us an upper limit for the dark matter particles
@@
which is in good agreement to the result obtained in (Ivanytskyi et al., 2020) ($m_{\chi}\leq 60\ \mathrm{GeV}$) in spite the fact that we are using a different EOS model for ordinary matter.
Notice that in the DM accretion rate estimate made in (Kouvaris, 2008), only the dark matter-nucleon interaction cross-section $\sigma_{\chi n}$ is taken into account. If a non zero self-interaction between dark matter particles is assumed, the accretion rate may change drastically (Güver et al., 2014; Nelson et al., 2019). This is because self-interactions between DM particles can dominate the capture process when the dark matter-nucleon cross-section $\sigma_{\chi n}$ is very small.
Moreover, other sources of DM accretion onto neutron stars should be considered. For example, in the central region of the Galaxy, clumps of DM could interact with the neutron star or its progenitor star leading to a large amount of accreted DM on the star. In (Deliyergiyev et al., 2024), without taking into account DM self-interaction, an accretion of $M_{DM}^{acc}\approx 7.5\times 10^{-4}\ \mathrm{M_{\odot}}$ is estimated for a typical neutron star inside a DM clump.
Finally, a significant amount of dark matter could be generated during the supernova explosion, with the majority remaining trapped within the neutron star (Nelson et al., 2019). Taking into account all these contributions and the uncertainties associated with each, one might arrive at a significantly higher estimate ($f_{\chi}\lesssim 10^{-2}$) of the DM fraction in a neutron star, with respect to the one ($f_{\chi}\sim 10^{-4}$) mentioned before and reported in Fig. (7).
The DM fraction values we consider in the present work (Fig. (4) and Fig. (5)) and commonly considered in the literature (e.g.(Ellis et al., 2018; Ivanytskyi et al., 2020; Grippa et al., 2024)) are thus compatible with the upper estimate ($f_{\chi}\lesssim 10^{-2}$) discussed above.
6 Stability of DANS
A solution the two-fluid TOV equations (2) and (2) represents an equilibrium configuration. In this section we present results regarding the stability with respect to small radial perturbation of DANS. We assume that these perturbations do not modify the $\beta$-equilibrium condition of ordinary matter.
In the one-fluid limit (stars made only of ordinary matter or only of dark matter) it is well known that the critical point is the maximum of the $M(\rho_{c,j})$ curve (Shapiro and Teukolsky, 1983). Using the EOSs considered in this work we get the following value for the central mass densities of the critical configuration
@@ 6 Stability of DANS
In the two-fluids system case the stability analysis can be done in two different ways. The first method, discussed in (Kain, 2020), requires us to solve a set of pulsation equations (eigenvalues problem) whose solutions give the squared radial oscillation frequencies $\omega^{2}_{n}(\rho_{cOM},\rho_{cDM})$ of the different modes. For any choice of the central densities of the two fluids $(\rho_{cOM},\rho_{cDM})$, if the squared radial oscillation frequency for the fundamental mode $\omega_{0}^{2}(\rho_{cOM},\rho_{cDM})$ is positive, then the corresponding solution (DANS) of the two fluids TOV equations is stable with respect to small radial perturbations while, if $\omega_{0}^{2}(\rho_{cOM},\rho_{cDM})<0$, the solution is unstable. The points in the two central density plane in which we have a transition from stable to unstable solutions are defined by the condition $\omega_{0}^{2}(\rho_{cOM},\rho_{cDM})=0$: this indicates that in the case of two-fluid systems we have a critical stability curve that corresponds to the geometric locus of points on the plane $(\rho_{cOM},\rho_{cDM})$ for which $\omega_{0}^{2}=0$. Determine the critical stability curve with this approach, requires us to take the derivative of the EOSs of the fluids which is a problem if the EOSs aren’t smooth enough (as in the case of the BL EOS, which is given in tabular form).
Since we are just interested in determining the stability region on the plane of the two central densities, a second method can be used to study the stability of two-fluids systems. This second approach, discussed by (Henriques et al., 1990) for Fermion-Boson stars, allows us to determine the critical stability curve without explicitly solving the pulsation equations. It can be shown that the critical curve is defined by the following set of equations
@@
where $M_{tot}$, $N_{i}$ ($i=OM,DM$) are the total gravitational mass and total fluid number of a static solution of the two-fluids TOV equations and $\vec{\sigma}$ is a vector in the parameter space (in this case the two central densities plane) that is simultaneously tangent to the level curves of $M_{tot}$ and $N_{i}$.
@@
It can be shown that to determine the critical stability curve, we can study the stationary point of $M_{tot}(\rho_{cOM},\rho_{cDM})$ along level curve of either $N_{OM}(\rho_{cOM},\rho_{cDM})$ or $N_{DM}(\rho_{cOM},\rho_{cDM})$ (see also (Valdez-Alvarado et al., 2013; Gleason et al., 2022)). As observed in (Henriques et al., 1990), the stability region must contain the stable structures in the one-fluid limit.
We apply the latter method in the $m_{\chi}=1\ \mathrm{GeV}$ case (see results in Fig.(10)) by studying stationary points of $M_{tot}(\rho_{cOM},\rho_{cDM})$ along $N_{DM}(\rho_{cOM},\rho_{cDM})$ level curve, in order to check that in our model we get a reasonably large stability region in the central densities plane.
@@
Our results are in agreement with those obtained in (Kain, 2021), and show us that we can have stable DANS with both a DM core or a DM halo (cfr central panel of Fig 2) and that exists a region inside the stability curve where DANS have an ordinary matter central density and/or a dark matter central density that exceed the critical value in the one-fluid limit.
Figure 10: Critical stability curve $\omega_{0}^{2}(\rho_{cOM},\rho_{cDM})=0$ for $m_{\chi}=1\ \mathrm{GeV}$ plotted as a contour line in the two central densities plane. Stable DANS configurations ($\omega_{0}^{2}>0$) are located in the lower left region of the two density plane, bounded by the black line. The different shades of color denote the function $M_{tot}(\rho_{cOM},\rho_{cDM})$.
We plot the $\Delta R=R_{OM}-R_{DM}=0$ contour line (white dashed-dotted line) in order to show that both stable DANS with a DM core and DANS with a DM halo are possible solutions to the two-fluid TOV equations (cfr. Fig.2).
7 Summary